OpenStudy (anonymous):

There are 9 true/false questions on a quiz. If twice as many of the statements are true as false,how many different sequences of true/fales answers are possible? There are 9 true/false questions on a quiz. If twice as many of the statements are true as false,how many different sequences of true/fales answers are possible? @Mathematics

OpenStudy (anonymous):

you have 6 true and 3 false right? so it is \[\dbinom{9}{3}=\dbinom{9}{6}=\frac{9\times 8\times 7}{3\times 2}\]